OpenStudy (doc.brown):

In empty space, an object appears 70cm away from a 9.8kg mass. What is its acceleration towards the mass?

OpenStudy (anonymous):

is there a time.. or any other info?

OpenStudy (anonymous):

is there any more info? anyways here are 4 equations that deal with acceleration: a = f/m and a = change in v/ change in time and s = s +vt+(1/2)at^2 and v^2 = v^2 + 2as

OpenStudy (compphysgeek):

Empty space means there are only those two masses. the mass that appears will only feel the gravitational force of the body M = 9.8 kg. Gravitational force is\[F = g \frac{m M}{r^2}\] M is the body's mass, r = 0.7 m is the distance between the objects, g is the gravitational constant. F is accelerating the object. \[ F = m a = g \frac{m M}{r^2}\] \[=> a = g \frac{M}{r^2}\]

OpenStudy (doc.brown):

Is this right?\[a=\frac{6.67\times 10^{-11}N\cdot m^2/kg^2\times 9.8kg}{(0.70m)^2}=1.3\times 10^{-9}m/s^2\]Is this right?

OpenStudy (alistairgarratt):

I would have though that you also have to think about the mass of the "object" in your question. Best wishes, Alistair

OpenStudy (doc.brown):

If the mass doubles, the force accelerating the particle also doubles. The acceleration stays the same. Can someone please confirm \(1.3\times 10^{-9}\)?

OpenStudy (compphysgeek):

I've also got \(a = 1.3 \times 10^{-9} ms^{-2}\)

OpenStudy (doc.brown):

Thank you kindly.